If you placed $2,000 into an investment account earning 3% simple interest, how many months does it take for you to have $2,025 in your account?
Answer #1
where,
P = principal
N = Number of years
R = Rate
Interest earned = $2025-2000
=$25
Let us calculate N
$25 =$2000*N*0.03
25 =60*N
N = 25/60
N = 0.416667
Number of months = 0.41667*12
=5 months

Please upvote.